OpenStudy (anonymous):

Solve the system. If there is no solution or if the system's equations are dependent, so state x + y + z = 4 x-y+3z= 10 2x + y + z = 1

OpenStudy (anonymous):

X = (-12/4) =-3 Y = (8/4) =2 Z = (20/4) = 5